What is the complete set of quantum numbers for the last entering electron of Nickel, Xenon and Bromine?


1
Expert's answer
2022-04-20T16:33:03-0400

Nickel:


Principal = 3 (because the electron is in energy level 3)

Angular = 2 (because the electron is in a d sublevel)

Magnetic = 0 (because the electron is in the middle orbital of the 5 orbitals in the 3d sublevel)

Spin = -1/2 (because the electron is down spin)


Xenon:

Principal = 5

Angular = 1

Magnetic = 1

Spin = -1/2


Bromine:

Principal = 4

Angular = 1

Magnetic = 0

Spin = -1/2